OpenStudy (anonymous):

Find the locations of the points with minimum distance to x^4yz^4 = 8√2 to the origin.

OpenStudy (anonymous):

you function would be: \(f(x,y,z)=x^2+y^2+z^2\) constraint: \(x^4yz^4 - 8√2=0\) Using Lagrange multipliers: \(df+\lambda dg=0\) You will get a system with 4 equations and four unknowns. Solve it and later check either it gives you mínimum or maximum

OpenStudy (anonymous):

Forgot to say what g is: g(x,y,z)=x^4yz^4 -8√2